Find the equilibrium quantity and equilibrium price for the commodity whose supply and demand functions are given.
​Supply: p=q^+20q
    
​Demand: p=-2q^+10q+11,400
The equilibrium quantity is q=
The equilibrium price is q=

Answer by stanbon(75887)

Find the equilibrium quantity and equilibrium price for the commodity whose supply and demand functions are given.
​Supply: p=q^+20q
​Demand: p=-2q^+10q+11,400
-------
Solve:: q^2+20q = -2q^2+10q+11400
3q^2 + 10q - 11400 = 0
(q-60)(3q+190) = 0
q = 60
Supply = q^2+20q = 60^2+20*60 = 4800
Demand = -2q^2+10q+11,400 = -2*60^2+10*60+11400 = 4800
-----------------------------------------
The equilibrium quantity is q=
The equilibrium price is q=
-----
Cheers,
Stan H.
-----------